A barometer containing mercury is owered in a vessel containing mecury until only 50 cm of  the tube is above the level of the mercury in the vessel. If the atmosphere pressure is 75 cm of mercury, what is the pressure at the top of the tube?

Grade:12

1 Answers

Rituraj Tiwari
3 years ago
Here, the mercury rises only 50 cm of the tube. and rest of the excess atmospheric pressure is being countered by the pressure at the top of the tube.

=> Atmospheric pressure of mercury = 75 cm

∴ Pressure(at top) + Pressure (due to 50cmHG) = atmospheric Pressure

Pressure(at top) = atmospheric Pressure - Pressure (due to 50cmHG)

= 75 - 50 cm of HG

= 25 cm of HG

= 25/75 * 10⁵ Pa

= 0.3333 * 10⁵ Pa

= 33.33 * 10³ Pa

= 33.33 K Pa

Thus, The pressure at the top of the tube is 33.33 K Pa.